Can you please help me understand how to solve
Y = -x2 + 2x + 4
x+y = 4
(that is -x squared)

Answer by ikleyn(52776)

.

From the second equation, you express  y = 4-x  and substitute it into the first equation, replacing "y" there


    4 - x = -x^2 + 2x + 4.


Now you just have one single quadratic equation.


Reduce it to the standard form quadratic equation, collecting all the terms in one side (on the left, in this case)


    x^2 - 3x = 0.


Solve by factoring


    x*(x-3) = 0.


The roots are x= 0  and  x= 3.


The solutions to the system are  (x,y) = (0,4)  and  (x,y) = (3,1).    ANSWER


You can check the answer by substituting the found solution values of x and y into the original system.

Solved and explained in all details.

Is everything clear to you ?

If you still have questions, do not hesitate to ask me.

-----------------

If you want to see many other similar solved problems, look into the lesson
    - Solving systems of algebraic equations of degree 2 and degree 1
in this site.